# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 4. Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
Let all the conditions in Theorem 4.1 be satisfied. There exists some constant $c=c_{p,q,h,\widehat{\lambda}}\in(0,+\infty)$ for $h\geq 3$, such that uniformly for partitions $I,J$ for $\{1,\cdots,h\}$ with $11\leq|I|,|J|\leq h-$ 1 and $I\neq J$ for $|I|=|J|=h-1$, we have that for large $N\leq\widetilde{N}$ and $\lambda=\frac{\widehat{\lambda}}{{N}}$,
(4.19) $\Big{\|}Q_{\lambda,N}^{I,J}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq c,$
(4.20) $\Big{\|}Q_{\lambda,N}^{*,I}\Big{\|}_{\ell^{q}\to\ell^{q}}$ $\leq cN^{\frac{1}{q}},\quad\Big{\|}Q_{\lambda,N}^{I,*}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq cN^{\frac{1}{p}},$
[…]
We postpone the proof of this proposition and apply it to completing the proof of Theorem 4.1 first. For $h\geq 3$, note that ${\mathbbm{E}}[\zeta^{I}]=\sigma_{\widetilde{N}}^{2}$ for $|I|=h-1$ and $|{\mathbbm{E}}[\zeta^{I}]|\leq c^{\prime}\sigma_{\widetilde{N}}^{3}=O((\log\widetilde{N})^{-{3/2}})$ for $|I|<h-1$. Then the above bounds for operators imply that for $2\leq i\leq r$,
(4.22) ${\mathbbm{E}}[\zeta^{I_{i}}]\Big{\|}P^{I_{i-1},I_{i}}_{\lambda}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq\mathbbm{1}_{\{|I_{i}|=h-1\}}\frac{c^{2}}{\log\big{(}\widehat{\lambda}\frac{\widetilde{N}}{N}\big{)}}+\mathbbm{1}_{\{|I_{i}|<h-1\}}cc^{\prime}\sigma_{\widetilde{N}}^{3}\leq\frac{c^{\prime\prime}}{\log\big{(}\widehat{\lambda}\frac{\widetilde{N}}{N}\big{)}},$
@@
(4.23) ${\mathbbm{E}}[\zeta^{I_{1}}]\Big{\|}P^{\ast,I_{1}}_{\lambda}\Big{\|}_{\ell^{q}\to\ell^{q}}\leq N^{\frac{1}{q}}\Big{(}\mathbbm{1}_{\{|I_{1}|=h-1\}}\frac{c}{\log\big{(}\widehat{\lambda}\frac{\widetilde{N}}{N}\big{)}}+\mathbbm{1}_{\{|I_{1}|<h-1\}}c^{\prime}\sigma_{\widetilde{N}}^{3}\Big{)}\leq\frac{c^{\prime\prime}}{\log(\widehat{\lambda}\frac{\widetilde{N}}{N})}N^{\frac{1}{q}}.$
For fixed $h$, the number of partitions $I_{i}$ is bounded by a constant $c_{h}$. Choose $\widehat{\lambda}$ large such that $c_{h}\frac{c^{\prime\prime}}{\log\widehat{\lambda}}\leq\frac{1}{2}$. Plug the bounds (4.22) and (4.23) into (4.18), and we have that for $N$ large enough,
(4.24) $\big{|}M_{N,\widetilde{N},h}^{\varphi,\psi}\big{|}\displaystyle\leq\frac{Ce^{2\widehat{\lambda}}\lVert\psi\rVert^{h}_{\infty}}{N^{\frac{h}{2}+1}}\big{\|}\varphi_{N}\big{\|}^{h}_{\ell^{p}}\|\mathbbm{1}_{B_{N}}\big{\rVert}^{h}_{\ell^{q}}N^{\frac{1}{p}+\frac{1}{q}}\sum_{r=1}^{\infty}\Big{(}\frac{c_{h}c^{\prime\prime}}{\log(\widehat{\lambda}\frac{\widetilde{N}}{N})}\Big{)}^{r}\!\!\leq\!\!\frac{C_{1}}{N^{\frac{h}{2}}\log(1+\frac{\widetilde{N}}{N})}\big{\|}\psi\big{\rVert}^{h}_{\infty}\big{\|}\varphi_{N}\big{\rVert}^{h}_{\ell^{p}}\big{\lVert}\mathbbm{1}_{B_{N}}\big{\rVert}^{h}_{\ell^{q}}.$
@@
(4.25) $Q_{\lambda,N}(\boldsymbol{x},\boldsymbol{y}):=\sum_{n=1}^{2N}e^{-\lambda n}\prod_{i=1}^{h}p_{n}(y_{i}-x_{i}),$
which is actually $Q_{\lambda,N}^{*,*}(\boldsymbol{x},\boldsymbol{y})$, a partition-free version of (4.14). Since $Q_{N}(\boldsymbol{x},\boldsymbol{y}):=Q_{0,N}(\boldsymbol{x},\boldsymbol{y})=\sup_{\lambda\geq 0}Q_{\lambda,N}(\boldsymbol{x},\boldsymbol{y})$, we only need to bound $Q_{N}(\boldsymbol{x},\boldsymbol{y})=\sum_{n=1}^{2N}\prod_{i=1}^{h}p_{n}(y_{i}-x_{i})$.
Note that ${\mathbf{E}}[|S_{1}|^{4}]<+\infty$ and ${\mathbf{E}}[(S_{1})^{3}]=0$. By [LL10, Theorem 2.3.10], there exists a constant $C\in(0,+\infty)$, such that uniformly in $x$ and $n$,
(4.26) $p_{n}(x)\leq g_{n}(x)+\frac{C}{(1+|x|^{2})\sqrt{n}},$
@@
(4.27) $\begin{split}Q_{N}(\boldsymbol{x},\boldsymbol{y})&\leq\sum\limits_{n=1}^{2N}\prod\limits_{i=1}^{h}\Big{(}g_{n}(y_{i}-x_{i})+\frac{C}{(1+|y_{i}-x_{i}|^{2})\sqrt{n}}\Big{)}\\ &=\sum\limits_{n=1}^{2N}\prod\limits_{i=1}^{h}g_{n}(y_{i}-x_{i})+\sum\limits_{n=1}^{2N}\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{2})\sqrt{n}}.\end{split}$
To bound the above quantity, we need [CSZ23, Lemma 6.7], which shows that there exists some constant $C\in(0,+\infty)$, such that uniformly in $x,y,N$ and $K\geq 3$,
(4.28) $\sum\limits_{i=1}^{2N}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\leq\begin{cases}\displaystyle\frac{C}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{|K|}{2}-1}},\quad&\text{for all}~{}\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})^{|K|},\\[16.0pt] \displaystyle\frac{C}{N^{\frac{|K|}{2}-1}}e^{-\frac{|\boldsymbol{y}-\boldsymbol{x}|^{2}}{CN}},\quad&\text{for}~{}|\boldsymbol{x}-\boldsymbol{y}|\geq\sqrt{N}.\end{cases}$
@@
(4.29) $\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})Q_{\lambda}^{I,J}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq c\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}.$
@@
(4.30) $\begin{split}&\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})Q_{\lambda}^{I,J}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})\frac{C}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{h}{2}-1}}\psi(\boldsymbol{y})\\ &+\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})\bigg{(}\frac{C}{(1+\sum_{i\in K}|y_{i}-x_{i}|^{2})^{\frac{h}{2}-1}}\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{2})}\bigg{)}\psi(\boldsymbol{y}).\end{split}$
@@
Fix $p,q>1$ with $\frac{1}{p}+\frac{1}{q}=1$ and an integer $h\geq 3$. Consider partitions $I,J$ with $11\leq|I|,|J|\leq h-$ 1, and $I\neq J$ if $|I|=|J|=h-1$. Let $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{q}(({\mathbbm{Z}})^{h}_{J})$. Then there exists $C=C_{p,q,h}<\infty$, independent of $f$ and $g$, such that
(4.31) $\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\varphi(\boldsymbol{x})\psi(\boldsymbol{y})}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{h}{2}-1}}\leq C\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}},$
@@
(4.32) $\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})\bigg{(}\frac{C}{(1+\sum_{i\in K}|y_{i}-x_{i}|^{2})^{\frac{h}{2}-1}}\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{2})}\bigg{)}\psi(\boldsymbol{y})\leq C\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}.$
@@
(4.33) $\sum_{x\in\mathbb{Z}}\frac{1}{(s+x^{2})^{r}}\leq\frac{C}{s^{r-1/2}},\quad\text{with a uniform $C>0$ for}~{}s>1,r>\frac{1}{2}.$
We first treat (4.31) with $|I|=|J|=h-1$. Assume that the index sets for renewal random walks are $I(1)=\{k,l\}$ and $J(1)=\{m,n\}$, with $\{k,l\}\neq\{m,n\}$. Choose $0<a<\frac{1}{2(p\vee q)}$. By Höder’s inequality, the left-hand side of (4.31) is bounded from above by
(4.34) $\begin{array}[]{l}\displaystyle\left(\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\varphi(\boldsymbol{x})^{p}}{(1+\sum_{i=1}^{h}|x_{i}-y_{i}|^{2})^{\frac{h}{2}-1}}\cdot\frac{(1+|x_{m}-x_{n}|^{2a})^{p}}{(1+|y_{k}-y_{l}|^{2a})^{p}}\right)^{1/p}\\[5.0pt] \displaystyle\qquad\times\left(\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\psi(\boldsymbol{y})^{q}}{(1+\sum_{i=1}^{h}|x_{i}-y_{i}|^{2})^{\frac{h}{2}-1}}\cdot\frac{(1+|y_{k}-y_{l}|^{2a})^{q}}{(1+|x_{m}-x_{n}|^{2a})^{q}}\right)^{1/q}.\end{array}$
@@
Proof for Case (i). For the first factor above, note that $x_{k}=x_{l}=y_{m}=y_{n}=0$ in $({\mathbbm{Z}})^{h}_{J}$. Use (4.33) to sum over the remaining $h-4$ variables $y_{j}$ with $j\neq k,l,m,n$. We get an upper bound
(4.35) $\begin{array}[]{l}\displaystyle C\Bigg{(}\sum_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I}}f(\boldsymbol{x})^{p}(1+|x_{m}-x_{n}|^{2a})^{p}\sum_{y_{k},y_{l}\in{\mathbbm{Z}}}\frac{1}{(1+x_{m}^{2}+x_{n}^{2}+y_{k}^{2}+y_{l}^{2})^{\frac{h}{2}-1-\frac{h-4}{2}}(1+|y_{k}-y_{l}|^{2a})^{p}}\Bigg{)}^{1/p}.\end{array}$
@@
(4.36) $\begin{array}[]{l}\displaystyle\sum_{\widetilde{y}_{1},\widetilde{y}_{2}}\frac{C}{(1+x_{m}^{2}+x_{n}^{2}+\widetilde{y}_{1}^{2}+\widetilde{y}_{2}^{2})(1+\widetilde{y}_{1}^{2a})^{p}}\displaystyle\leq\sum_{\widetilde{y}_{1}}\frac{C}{(1+x_{m}^{2}+x_{n}^{2}+\widetilde{y}_{1}^{2})^{1/2}(1+\widetilde{y}_{1}^{2a})^{p}}\leq\frac{C}{(1+x_{m}^{2}+x_{n}^{2})^{ap}},\end{array}$
where the last inequality is achieved by summing over $|\widetilde{y}_{1}|\leq\sqrt{x_{m}^{2}+x_{n}^{2}}$ and $|\widetilde{y}_{1}|>\sqrt{x_{m}^{2}+x_{n}^{2}}$ separately. Then $(1+|x_{m}-x_{n}|^{2a})^{p}$ and (4.36) cancel out, and thus the first factor of (4.34) is upper bounded by $C\lVert\varphi\rVert_{\ell^{p}}$. The second factor of (4.34) can similarly bounded by $C\|\psi\|_{\ell^{q}}$.
Proof for Case (ii). For the first factor in (4.34), note that $x_{m}=x_{l}=y_{m}=y_{n}=0$. Sum over remaining $h-3$ variables $y_{j}$ with $j\neq l,m,n$. Similarly, we get an upper bound
(4.37) $\begin{array}[]{l}\displaystyle\quad C\Bigg{(}\sum_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I}}\varphi(\boldsymbol{x})^{p}(1+|x_{n}|^{2a})^{p}\sum_{y_{l}}\frac{1}{(1+x_{n}^{2}+y_{l}^{2})^{1/2}(1+y_{l}^{2a})^{p}}\Bigg{)}^{1/p}\leq C\|\varphi\|_{\ell^{p}},\\[5.0pt] \end{array}$

### v3

@@ 4 Moment estimates for ${\mathcal{Z}}_{0,N}^{\beta_{N}}(\varphi,\psi)$
Let all the conditions in Theorem 4.1 be satisfied. There exists some constant $c=c_{p,q,h,\widehat{\lambda}}\in(0,+\infty)$ for $h\geq 3$, such that uniformly for partitions $I,J$ for $\{1,\cdots,h\}$ with $1\leq|I|,|J|\leq h-1$ and $I\neq J$ for |I|=|J|=h1|I|=|J|=h-1, we have that for large $N\leq\widetilde{N}$ and $\lambda=\frac{\widehat{\lambda}}{{N}}$,
$\Big\|Q_{\lambda,N}^{I,J}\Big\|_{\ell^{q}\to\ell^{q}}\leq c,$ (4.19)
$\Big\|Q_{\lambda,N}^{*,I}\Big\|_{\ell^{q}\to\ell^{q}}$ $\leq cN^{\frac{1}{q}},\quad\Big\|Q_{\lambda,N}^{I,*}\Big\|_{\ell^{q}\to\ell^{q}}\leq cN^{\frac{1}{p}},$ (4.20)
$\text{and in particular, for}~|I|=h-1$ $,\quad\Big\|U_{\lambda,N,\widetilde{N}}^{J}\Big\|_{\ell^{q}\to\ell^{q}}\leq\frac{c}{\log\big(\widehat{\lambda}\frac{\widetilde{N}}{N}\big)\sigma_{\widetilde{N}}^{2}}.$ (4.21)
We postpone the proof of this proposition and apply it to completing the proof of Theorem 4.1 first. For $h\geq 3$, note that ${\mathbbm{E}}[\zeta^{I}]=\sigma_{\widetilde{N}}^{2}$ for |I|=h1|I|=h-1 and $|{\mathbbm{E}}[\zeta^{I}]|\leq c^{\prime}\sigma_{\widetilde{N}}^{3}=O((\log\widetilde{N})^{-{3/2}})$ for |I|<h1|I|<h-1. Then the above bounds for operators imply that for $2\leq i\leq r$,
${\mathbbm{E}}[\zeta^{I_{i}}]\Big\|P^{I_{i-1},I_{i}}_{\lambda}\Big\|_{\ell^{q}\to\ell^{q}}\leq\mathbbm{1}_{\{|I_{i}|=h-1\}}\frac{c^{2}}{\log\big(\widehat{\lambda}\frac{\widetilde{N}}{N}\big)}+\mathbbm{1}_{\{|I_{i}|<h-1\}}cc^{\prime}\sigma_{\widetilde{N}}^{3}\leq\frac{c^{\prime\prime}}{\log\big(\widehat{\lambda}\frac{\widetilde{N}}{N}\big)},$ (4.22)
@@
${\mathbbm{E}}[\zeta^{I_{1}}]\Big\|P^{\ast,I_{1}}_{\lambda}\Big\|_{\ell^{q}\to\ell^{q}}\leq N^{\frac{1}{q}}\Big(\mathbbm{1}_{\{|I_{1}|=h-1\}}\frac{c}{\log\big(\widehat{\lambda}\frac{\widetilde{N}}{N}\big)}+\mathbbm{1}_{\{|I_{1}|<h-1\}}c^{\prime}\sigma_{\widetilde{N}}^{3}\Big)\leq\frac{c^{\prime\prime}}{\log(\widehat{\lambda}\frac{\widetilde{N}}{N})}N^{\frac{1}{q}}.$ (4.23)
For fixed $h$, the number of partitions $I_{i}$ is bounded by a constant $c_{h}$. Choose $\widehat{\lambda}$ large enough such that $c_{h}\frac{c^{\prime\prime}}{\log\widehat{\lambda}}\leq\frac{1}{2}$. Plug the bounds (4.22) and (4.23) into (4.18), and we have that for $N$ large enough,
$\begin{split}\big|M_{N,\widetilde{N},h}^{\varphi,\psi}\big|\displaystyle\leq&\frac{Ce^{2\widehat{\lambda}}\lVert\psi\rVert^{h}_{\infty}}{N^{\frac{h}{2}+1}}\big\|\varphi_{N}\big\|^{h}_{\ell^{p}}\|\mathbbm{1}_{B_{N}}\big\rVert^{h}_{\ell^{q}}N^{\frac{1}{p}+\frac{1}{q}}\sum_{r=1}^{\infty}\Big(\frac{c_{h}c^{\prime\prime}}{\log(\widehat{\lambda}\frac{\widetilde{N}}{N})}\Big)^{r}\\ \leq&\frac{C_{1}}{N^{\frac{h}{2}}\log(1+\frac{\widetilde{N}}{N})}\big\|\psi\big\rVert^{h}_{\infty}\big\|\varphi_{N}\big\rVert^{h}_{\ell^{p}}\big\lVert\mathbbm{1}_{B_{N}}\big\rVert^{h}_{\ell^{q}}.\end{split}$ (4.24)
@@
$Q_{\lambda,N}(\boldsymbol{x},\boldsymbol{y}):=\sum_{n=1}^{2N}e^{-\lambda n}\prod_{i=1}^{h}p_{n}(y_{i}-x_{i}),$ (4.25)
which is actually $Q_{\lambda,N}^{*,*}(\boldsymbol{x},\boldsymbol{y})$, a partition-free version of (4.14). Since $Q_{N}(\boldsymbol{x},\boldsymbol{y}):=Q_{0,N}(\boldsymbol{x},\boldsymbol{y})=\sup_{\lambda\geq 0}Q_{\lambda,N}(\boldsymbol{x},\boldsymbol{y})$, it suffices to bound $Q_{N}(\boldsymbol{x},\boldsymbol{y})=\sum_{n=1}^{2N}\prod_{i=1}^{h}p_{n}(y_{i}-x_{i})$.
By Theorem B.2, there exist constants $\gamma,C>0$, such that uniformly in $x$ and $n$,
$p_{n}(x)\leq g_{n}(x)+\frac{C}{(1+|x|^{1+\gamma})\sqrt{n}},$ (4.26)
@@
$\begin{split}Q_{N}(\boldsymbol{x},\boldsymbol{y})&\leq\sum\limits_{n=1}^{2N}\prod\limits_{i=1}^{h}\Big(g_{n}(y_{i}-x_{i})+\frac{C}{(1+|y_{i}-x_{i}|^{1+\gamma})\sqrt{n}}\Big)\\ &=\sum\limits_{n=1}^{2N}\prod\limits_{i=1}^{h}g_{n}(y_{i}-x_{i})+\sum\limits_{n=1}^{2N}\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{1+\gamma})\sqrt{n}}.\end{split}$ (4.27)
To bound the above quantity, we need [15, Lemma 6.7], which shows that there exists some constant $C\in(0,+\infty)$, such that uniformly in $x,y,N$ and $K\geq 3$,
$\sum\limits_{n=1}^{2N}\prod\limits_{i\in K}g_{n}(y_{i}-x_{i})\leq\begin{cases}\displaystyle\frac{C}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{K}{2}-1}},\quad&\text{for all}~\boldsymbol{x},\boldsymbol{y}\in({\mathbbm{Z}})^{K},\\[16.0pt] \displaystyle\frac{C}{N^{\frac{K}{2}-1}}e^{-\frac{|\boldsymbol{y}-\boldsymbol{x}|^{2}}{CN}},\quad&\text{for}~|\boldsymbol{x}-\boldsymbol{y}|\geq\sqrt{N}.\end{cases}$ (4.28)
@@
$\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})Q_{\lambda}^{I,J}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq c\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}.$ (4.29)
@@
$\begin{split}&\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})Q_{\lambda}^{I,J}(\boldsymbol{x},\boldsymbol{y})\psi(\boldsymbol{y})\leq\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})\frac{C}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{h}{2}-1}}\psi(\boldsymbol{y})\\ &+\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\varphi(\boldsymbol{x})\bigg(\frac{C}{(1+\sum_{i\in K}|y_{i}-x_{i}|^{2})^{\frac{h}{2}-1}}\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{1+\gamma})}\bigg)\psi(\boldsymbol{y}).\end{split}$ (4.30)
@@
Fix $p,q>1$ with $\frac{1}{p}+\frac{1}{q}=1$ and an integer $h\geq 3$. Consider partitions $I,J$ with $1\leq|I|,|J|\leq h-1$, and $I\neq J$ if |I|=|J|=h1|I|=|J|=h-1. Let $\varphi\in\ell^{p}(({\mathbbm{Z}})^{h}_{I})$ and $\psi\in\ell^{q}(({\mathbbm{Z}})^{h}_{J})$. Then there exists $C=C_{p,q,h}<\infty$, independent of $f$ and $g$, such that
$\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\varphi(\boldsymbol{x})\psi(\boldsymbol{y})}{(1+|\boldsymbol{x}-\boldsymbol{y}|^{2})^{\frac{h}{2}-1}}\leq C\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}},$ (4.31)
@@
$\begin{split}\sum\limits_{K\subsetneqq\{1,\cdots,h\}}\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}&\varphi(\boldsymbol{x})\bigg(\frac{C}{(1+\sum_{i\in K}|y_{i}-x_{i}|^{2})^{\frac{h}{2}-1}}\prod\limits_{j\notin K}\frac{C}{(1+|y_{j}-x_{j}|^{1+\gamma})}\bigg)\psi(\boldsymbol{y})\\ \leq&C\|\varphi\|_{\ell^{p}}\|\psi\|_{\ell^{q}}\end{split}$ (4.32)
@@
$\sum_{x\in\mathbb{Z}}\frac{1}{(s+x^{2})^{r}}\leq\frac{C}{s^{r-1/2}},\quad\text{with a uniform $C>0$ for}~s>1,r>\frac{1}{2}.$ (4.33)
We first treat (4.31) with |I|=|J|=h1|I|=|J|=h-1. Assume that the index sets for renewal random walks are $I(1)=\{k,l\}$ and $J(1)=\{m,n\}$, with $\{k,l\}\neq\{m,n\}$. Choose $0<a<\frac{1}{2(p\vee q)}$. By Hölder’s inequality, the left-hand side of (4.31) is bounded from above by
$\begin{array}[]{l}\displaystyle\left(\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\varphi(\boldsymbol{x})^{p}}{(1+\sum_{i=1}^{h}|x_{i}-y_{i}|^{2})^{\frac{h}{2}-1}}\cdot\frac{(1+|x_{m}-x_{n}|^{2a})^{p}}{(1+|y_{k}-y_{l}|^{2a})^{p}}\right)^{1/p}\\[5.0pt] \displaystyle\qquad\times\left(\sum\limits_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I},\boldsymbol{y}\in({\mathbbm{Z}})^{h}_{J}}\frac{\psi(\boldsymbol{y})^{q}}{(1+\sum_{i=1}^{h}|x_{i}-y_{i}|^{2})^{\frac{h}{2}-1}}\cdot\frac{(1+|y_{k}-y_{l}|^{2a})^{q}}{(1+|x_{m}-x_{n}|^{2a})^{q}}\right)^{1/q}.\end{array}$ (4.34)
@@
Proof for Case (i). For the first factor above, note that $x_{k}=x_{l}=y_{m}=y_{n}=0$ in $({\mathbbm{Z}})^{h}_{J}$. Use (4.33) to sum over the remaining h4h-4 variables $y_{j}$ with $j\neq k,l,m,n$. It is bounded above by a constant timing
$\begin{array}[]{l}\displaystyle\Bigg(\sum_{\boldsymbol{x}\in({\mathbbm{Z}})^{h}_{I}}\varphi(\boldsymbol{x})^{p}(1+|x_{m}-x_{n}|^{2a})^{p}\!\!\sum\limits_{y_{k},y_{l}\in{\mathbbm{Z}}}\frac{1}{(1+x_{m}^{2}+x_{n}^{2}+y_{k}^{2}+y_{l}^{2})(1+|y_{k}-y_{l}|^{2a})^{p}}\Bigg)^{1/p}.\end{array}$ (4.35)
@@
$\begin{split}\displaystyle&\sum\limits_{\widetilde{y}_{1},\widetilde{y}_{2}}\frac{C}{(1+x_{m}^{2}+x_{n}^{2}+\widetilde{y}_{1}^{2}+\widetilde{y}_{2}^{2})(1+\widetilde{y}_{1}^{2a})^{p}}\\ \displaystyle\leq&\sum\limits_{\widetilde{y}_{1}}\frac{C}{(1+x_{m}^{2}+x_{n}^{2}+\widetilde{y}_{1}^{2})^{1/2}(1+\widetilde{y}_{1}^{2a})^{p}}\leq\frac{C}{(1+x_{m}^{2}+x_{n}^{2})^{ap}},\end{split}$ (4.36)
where the last inequality is achieved by summing over $|\widetilde{y}_{1}|\leq\sqrt{x_{m}^{2}+x_{n}^{2}}$ and $|\widetilde{y}_{1}|>\sqrt{x_{m}^{2}+x_{n}^{2}}$ separately. Then $(1+|x_{m}-x_{n}|^{2a})^{p}$ and (LABEL:HLS3) cancel out, and thus the first factor of (4.34) is upper bounded by $C\lVert\varphi\rVert_{\ell^{p}}$. The second factor of (4.34) can similarly be bounded by $C\|\psi\|_{\ell^{q}}$.
Proof for Case (ii). For the first factor in (4.34), note that $x_{m}=x_{l}=y_{m}=y_{n}=0$. Sum over the remaining h3h-3 variables $y_{j}$ with $j\neq l,m,n$. Similarly, we get an upper bound
[…]